# Recursions, Trains, Trees, and Combinatorial Rod Set Algebra

### v1

@@ 11 Fibonacci and Lucas 2-Expansions
Since odd sign swap 20 preserves the magnitudes of the net train counts, this argument works for ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$ as well. ∎
@@
We need prove only existence; Theorem 52 guarantees uniqueness.
It suffices to find the expansions for the positive Lucas rod sets $1^{s},2^{t}]$. Then the invariance of the absolute values of train counts under odd sign swap 20 implies the expansions exist for ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$.
@@
Every other count is a multiple of $3$, as predicted by Lemma 54.
@@
    To find the expansion to shape $\langle a,a+2\rangle$ when $s=1$ or $a$ is even, we note that Lemma 54 implies those conditions hold just when $t\operatorname{F}(a-1)/s$ is an integer. Then we start with any rod set in Equation 11.2 and add $t\operatorname{F}(a-1)/s$ rod/antirod pairs of length $a$, expand the new antirods, and simplify. For example
@@
    To find more expansions with shape lengths that differ by $d=4$, we start with the rod set $4^{161},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{8}}^{16}]$ and repeatedly expand the smallest rods. Corollary 22 implies the expansions we create are also expansions of $1^{3},2^{2}]$ by finite $Q$. Below are the first few such expansions in our example:
@@
Theorem 55 proves the existence of the expansions, which together with Equation 10.4 in the 2-shape Theorem 52 imply the needed divisibility relationships between net train counts:
@@
    The expansions to rod sets with lengths $\langle a,a+2\rangle$ for odd $a$ imply the divisibilities $\operatorname{F}(1)=s\mid\operatorname{F}(a)$ for odd $a$. This is not a surprise, since we needed Lemma 54 to find those expansions.
@@
It’s well known that the traditionally indexed Fibonacci numbers are a strong divisibility sequence. It’s somewhat less well known that the Lucas sequences, as we have defined them are also strong divisibility sequences [1][5][6]. This result implies that the expansions we found in Theorem 55 lead to all possible two-term linear recursions involving Lucas sequences. In particular, we have found all such Fibonacci recursions.
@@ 12 Borwein trinomials
We write down the period $6$ or period $3$ net train counts for each rod set. Since the counts are all $\pm 1$ or 0 Theorem 52 says the expansions correspond to a scaling that matches 1’s or 1-1s at (ba1,b1)(b-a-1,b-1) with $\alpha=1$ or one that matches 1 and 1-1 with $\alpha=-1$. ∎
@@ 13 $2$-Expandability when $\max R>2$
So far we have applied Theorem 52 when $R$ contains only rods of lengths 1 and 2. We end this section (and the paper) with some examples and conjectures on $2$-expansions of rod sets $R$ with $\max R>2$.
@@
  • There is more to discover on the connections between expansions of $R$ and the structure of the net train count sequence $\operatorname{F}(n,R)$. Theorems 50 and 52 settle those connections for expansions to shapes of cardinality 1 and 2. What are the corresponding theorems for larger shapes?

### v2

@@ 11 Fibonacci and Lucas 2-Expansions
Since odd sign swap 5.4 preserves the magnitudes of the net train counts, this argument works for ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$ as well. ∎
@@
We need prove only existence; Theorem 10.6 guarantees uniqueness.
It suffices to find the expansions for the positive Lucas rod sets $1^{s},2^{t}]$. Then the invariance of the absolute values of train counts under odd sign swap 5.4 implies the expansions exist for ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{1}}^{s},2^{t}]$.
@@
Every other count is a multiple of $3$, as predicted by Lemma 11.2.
@@
    To find the expansion to shape $\langle a,a+2\rangle$ when $s=1$ or $a$ is even, we note that Lemma 11.2 implies those conditions hold just when $t\operatorname{F}(a-1)/s$ is an integer. Then we start with any rod set in Equation 11.2 and add $t\operatorname{F}(a-1)/s$ rod/antirod pairs of length $a$, expand the new antirods, and simplify. For example
@@
    To find more expansions with shape lengths that differ by $d=4$, we start with the rod set $4^{161},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{8}}^{16}]$ and repeatedly expand the smallest rods. Corollary 5.6 implies the expansions we create are also expansions of $1^{3},2^{2}]$ by finite $Q$. Below are the first few such expansions in our example:
@@
Theorem 11.3 proves the existence of the expansions, which together with Equation 10.4 in the 2-shape Theorem 10.6 imply the needed divisibility relationships between net train counts:
@@
    The expansions to rod sets with lengths $\langle a,a+2\rangle$ for odd $a$ imply the divisibilities $\operatorname{F}(1)=s\mid\operatorname{F}(a)$ for odd $a$. This is not a surprise, since we needed Lemma 11.2 to find those expansions.
@@
It’s well known that the traditionally indexed Fibonacci numbers are a strong divisibility sequence. It’s somewhat less well known that the Lucas sequences, as we have defined them are also strong divisibility sequences [1][6][7]. This result implies that the expansions we found in Theorem 11.3 lead to all possible two-term linear recursions involving Lucas sequences. In particular, we have found all such Fibonacci recursions.
@@ 12 Borwein trinomials
We write down the period $6$ or period $3$ net train counts for each rod set. Since the counts are all $\pm 1$ or 0 Theorem 10.6 says the expansions correspond to a scaling that matches 1’s or 1-1s at (ba1,b1)(b-a-1,b-1) with $\alpha=1$ or one that matches 1 and 1-1 with $\alpha=-1$. ∎
@@ 13 $2$-Expandability when $\max R>2$
So far we have applied Theorem 10.6 when $R$ contains only rods of lengths 1 and 2. We end this section (and the paper) with some examples and conjectures on $2$-expansions of rod sets $R$ with $\max R>2$.
@@
  • There is more to discover on the connections between expansions of $R$ and the structure of the net train count sequence $\operatorname{F}(n,R)$. Theorems 10.4 and 10.6 settle those connections for expansions to shapes of cardinality 1 and 2. What are the corresponding theorems for larger shapes?